OpenStudy (anonymous):

the area of a trapezoid is 184 square the height is 16cm and the length of one is the parallel side is 5cm find the length of the two parallel sides

OpenStudy (gudden):

@tempera85 ... area of trapezoid = 1/2 ( sum of parallel sides ) height and area is given to be 184 and height = 16 and length of one parallel side is 5 so let the other side be x using these values in formula we get 184 = 1/2 ( 16) ( sum of parallel sides ) on multiplying the equation with 2 we get 368 = 16 (sum of parallel sides) on dividing both sides by 16 we get 23 = sum of parallel sides